OpenStudy (perii224):

PLEASE HELP :( WILL MEDAL Select the equation in which the graph of the line has a negative slope, and the y-intercept equals 4/3 5x + 3y = 4 5x - 3y = –4 4x + 16y = –20 6x + 2y = –8 Select the equation that contains the point (4, -8), and in which the graph of the line has a positive slope. y – 8 = –4(x + 4) y – 8 = 4(x + 4) y + 8 = +4(x – 4) y + 8 = –4(x – 4)

OpenStudy (princessstyles-horan):

Okay well first you have to put the equations into slope-intercept form

OpenStudy (perii224):

Okay

OpenStudy (princessstyles-horan):

@perii224, then plug in the numbers

OpenStudy (perii224):

Okay, do you know how?

OpenStudy (princessstyles-horan):

Yes I'm in 8th, but I can help one second

OpenStudy (princessstyles-horan):

Okay so subtract 5x from both sides

OpenStudy (perii224):

Okay, thanks :)

OpenStudy (princessstyles-horan):

so it becomes 3y = -5x + 4

OpenStudy (perii224):

Oh okay, but that's not one of the options

OpenStudy (princessstyles-horan):

Then that one is not it. Move on the next one.

OpenStudy (perii224):

So you dont know what the first one is?

OpenStudy (princessstyles-horan):

the first one does have a negative slope, but does 4 = 4/3?

OpenStudy (perii224):

I guess

OpenStudy (princessstyles-horan):

No it doesn't

OpenStudy (perii224):

oh

OpenStudy (princessstyles-horan):

So next one...what grade are you in?

OpenStudy (perii224):

I'm in 9th

OpenStudy (princessstyles-horan):

oh okay....Hmmmm did you learn point-slope, and slope-intercept forms last year?

OpenStudy (perii224):

I dont remember xD

OpenStudy (princessstyles-horan):

oh okay

OpenStudy (perii224):

Do you know the second one?

OpenStudy (princessstyles-horan):

one sec

OpenStudy (perii224):

ok

OpenStudy (princessstyles-horan):

Okay so you have 5x - 3y = -4 subtract 5x from both sides -3y = -5x - 4 can't have a negative 3 so divide both sides by -3 to get y by itself y = 4/3 x - 4

OpenStudy (perii224):

Oh okay

OpenStudy (princessstyles-horan):

does that fit the description?

OpenStudy (perii224):

I guess, but what is the final answer?

OpenStudy (princessstyles-horan):

No it doesn't. keep doing that for all of the problems

OpenStudy (perii224):

Oh okay

OpenStudy (princessstyles-horan):

is this starting to make sense?

OpenStudy (perii224):

I guess

OpenStudy (princessstyles-horan):

that's not assuring